# SLAM; Point Cloud Registration; LiDAR; Multi-robot SLAM

### v2

@@ 3.2 Multi-robot map merging
Here we present the results of the map merging. Given the distributed approach that we are following, we use the term "map merging" because it integrates the loop closures of each robot altogether with the actual merging across the maps built by each robot. We test the map merging for each robot pair, resulting in 6 different configurations, obtained by varying the robots involved in the pair. We also analyze the impact of using PCM filtering and the use of all keyframes, i.e., those considered by the usual heuristics creating a keyframe every 0.5m motion by a robot, compared with using only tunnel-filtered keyframes. In tables 2 and 3, we show how our proposed filtering process, which we refer to as ”tunnel-filtered KF“ influences the percentage of outliers present in the final graphs. We can clearly see that our approach eliminates a large percentage of outliers, both wrong loops added because of a wrong place recognition detection, and wrong loops added because the place was correctly recognized but the point cloud registration algorithm failed. We consider the PCR failed if the translation error exceeds 1 meter or the rotation error exceeds 15 degrees. Moreover, in table 4 we show the impact of the reduction in the loop outliers and consider SLAM successful if the overall structure of the final optimized map is consistent with the expected map. Further visual information is provided in Appendix A and Appendix B.
@@
Table 4: “All KFs” shows the SLAM successes when all keyframes were used for loop detection, “Tunnel-filtered KFs” shows the SLAM success when filtering out tunnel-like point clouds from the loop detection. We define ”success“ if all the robots have a ratio between max ATE and path length less then 1%
@@
An interesting fact that can be noticed is that un-informative keyframes influence both the number of outliers in the Place Recognition task, due to global perceptual aliasing, and in the Point Cloud registration task, due to the low geometric structure that makes the alignment fail due to perceptual aliasing in local point cloud patches. By applying the keyframe filtering we retain only the point clouds that are both globally distinct from others and also present enough local distinctiveness to allow the correct estimation of the alignment.
@@
An especially noteworthy observation emerges from the configuration denoted as (0,2) with PCM filtering in Table 4. In this case, the SLAM pipeline fails when tunnel filtering is applied but succeeds when all keyframes are retained. This counterintuitive result underscores a critical limitation of our current keyframe filtering strategy: while intended to improve efficiency and robustness, the filtering method may inadvertently discard keyframes that contain essential spatial information necessary for accurate loop detection. This finding suggests a clear need for more refined and adaptive keyframe selection techniques, capable of preserving structurally significant frames, while discarding keyframes that are not useful for the loop detection.

### v3

@@ 3.2 Multi-robot map merging
Here we present the results of the map merging. Given the distributed approach that we are following, we use the term "map merging" because it integrates the loop detections of each robot altogether with the actual merging across the maps built by each robot. We test the map merging for each robot pair, resulting in 6 different configurations, obtained by varying the robots involved in the pair. We also analyze the impact of using PCM filtering and the use of all keyframes, i.e., those considered by the usual heuristics creating a keyframe every 0.5m motion by a robot, compared with using only tunnel-filtered keyframes. In tables 2 and 3, we show how our proposed filtering process, which we refer to as ”tunnel-filtered KF“ influences the percentage of outliers present in the final graphs. We can clearly see that our approach eliminates a large percentage of outliers, both wrong loops added because of a wrong place recognition detection, and wrong loops added because the place was correctly recognized but the point cloud registration algorithm failed. We consider the PCR failed if the translation error exceeds 1 meter or the rotation error exceeds 15 degrees. Moreover, in table 4 we show the impact of the reduction in the loop outliers and consider SLAM successful if the overall structure of the final optimized map is consistent with the expected map. Further visual information is provided in Appendix A and Appendix B.
@@
Table 4: Top rows without PCM filtering, bottom ones with PCM filtering. “All KFs” shows the SLAM successes when all keyframes were used for loop detection, “Tunnel-filtered KFs” shows the SLAM successes when filtering out tunnel-like point clouds from the loop detection. We define ”success“ if all the robots have a ratio between max ATE and path length less then 1%
@@
An interesting fact that can be noticed is that un-informative keyframes influence both the number of outliers in the Place Recognition task, due to global perceptual aliasing, and in the Point Cloud Registration task, due to the low geometric structure that makes the alignment fail due to perceptual aliasing in local point cloud patches. By applying our keyframe tunnel-filtering we retain only the point clouds that are both globally distinct from others and also present enough local distinctiveness to allow the correct estimation of the alignment.
@@
An especially noteworthy observation emerges from the configuration denoted as (0,2) with PCM filtering in Table 4. In this case, the SLAM pipeline fails when tunnel-filtering is applied but succeeds when all keyframes are retained. This counterintuitive result underscores a critical limitation of our current keyframe filtering strategy: while intended to improve efficiency and robustness, the filtering method may inadvertently discard keyframes that contain essential spatial information necessary for accurate loop detection. This finding suggests a clear need for more refined and adaptive keyframe selection techniques, capable of preserving structurally significant frames, while discarding keyframes that are not useful for the loop detection.